# TUTORIAL ARTICLE Contraction Theory for Nonlinear Stability Analysis and Learning-based Control: A Tutorial Overview

### v7

@@ 5.3 Motivation for Neural Contraction Metrics
The CV-STEM schemes in Theorems 4.24.5, and 4.6 permit the construction of optimal contraction metrics via convex optimization for synthesizing optimal, provably stable, and robust feedback control and state estimation. It is also shown in Theorems 5.2 and 5.3 that these contraction metrics are useful for obtaining stability guarantees for machine learning-based and data-driven control techniques while retaining the CV-STEM-type optimality due to the tracking error bounds (5.18) and (5.22), which are comparable to those in Theorems 4.1 and 4.4. However, the CV-STEM requires that a nonlinear system of equations or an optimization problem be solved at each time instant, which is not suitable for systems with limited online computational capacity.
@@
The Neural Contraction Metric (NCM) and its extensions Tsukamoto and Chung [2021], Tsukamoto et al. [2021, 2021], Tsukamoto and Chung [2021] have been developed to address such a computational issue by modeling the CV-STEM optimization scheme using a DNN Hornik et al. [1989], Cybenko [1989], Funahashi [1989], making it implementable in real-time. We will see in the subsequent sections that Theorems 5.2 and 5.3 provide one formal approach to prove robustness and stability properties of the NCM methodologies Tsukamoto et al. [2021].
6 Learning-based Robust Control and Estimation
@@
with $\sup_{x,t}\|d_{e0}\|=\bar{d}_{e0}\in\mathbb{R}_{\geq 0}$, $\sup_{x,t}\|d_{e1}\|=\bar{d}_{e1}\in\mathbb{R}_{\geq 0}$, $\sup_{x,t}\|G_{e0}\|_{F}=\bar{g}_{e0}\in\mathbb{R}_{\geq 0}$, and $\sup_{x,t}\|G_{e1}\|_{F}=\bar{g}_{e1}\in\mathbb{R}_{\geq 0}$. The following definitions of a DNN Hornik et al. [1989], Cybenko [1989], Funahashi [1989] and NCM Tsukamoto and Chung [2021], Tsukamoto et al. [2021, 2021] will be utilized extensively hereinafter.
@@
A neural network is a nonlinear mathematical model to approximately represent training samples $\{(x_{i},y_{i})\}_{i=1}^{N}$ of $y=\phi(x)$ by optimally tuning its hyperparameters $W_{\ell}$, and is given as follows:
@@
A Neural Contraction Metric (NCM) is a DNN model of a contraction metric given in Theorem 2.1. Its training data could be sampled by solving the CV-STEM presented in Theorem 4.2 for control (or Theorem 4.6 for differential feedback control) and Theorem 4.5 for estimation. The NCM framework is summarized in Fig. 3.
@@
Although a stochastic version of the NCM is called a Neural Stochastic Contraction Metric (NSCM) in Tsukamoto et al. [2021], we also denote it as an NCM in this paper for simplicity of presentation.
@@
6.1 Stability of NCM-based Control and Estimation
@@
Let $\mathcal{M}$ define the NCM in Definition 6.2, and let $u^{*}=u_{d}-R^{-1}B^{\top}M(x-x_{d})$ for $M$ and $R$ given in Theorem 4.2. Suppose that the systems (6.1) and (6.2) are controlled by $u_{L}$, which computes the CV-STEM controller $u^{*}$ of Theorem 4.2 replacing $M$ by $\mathcal{M}$, i.e.,
@@
for $\mathtt{e}=x-x_{d}$. Note that we could use the differential feedback control $u=u_{d}+\int^{1}_{0}kd\mu$ of Theorem 4.6 for $u^{*}$ and $u_{L}$. Define $\Delta_{L}$ of Theorems 5.2 and 5.3 as follows:
@@
where we use $q(0,t)=\xi_{0}(t)=x_{d}(t)$, $q(1,t)=\xi_{1}(t)=x(t)$, and $\varpi=[x^{\top},x_{d}^{\top},u_{d}^{\top}]^{\top}$ in the learning-based control formulation (5.1) and (5.2). If the NCM is learned to satisfy the learning error assumptions of Theorems 5.2 and 5.3 for $\Delta_{L}$ given by (6.8), then (5.18), (5.22), and (5.23) of Theorems 5.2 and 5.3 hold as follows:
@@
where $\mathtt{e}=x-x_{d}$, $C_{C}=\bar{g}_{c}^{2}({2}{\alpha_{G}}^{-1}+1)+\epsilon_{\ell 0}^{2}\alpha_{d}^{-1}$, the disturbance upper bounds $\bar{d}_{c}$ and $\bar{g}_{c}$ are given in (6.1) and (6.2), respectively, and the other variables are defined in Theorems 4.25.2 and 5.3.
@@
Let g of the virtual systems given in Theorems 5.2 and 5.3 be $\textsl{g}=\zeta$, where $\zeta$ is given by (3.19) (i.e., $\zeta=(A-BR^{-1}B^{\top}M)(q-x_{d})+\dot{x}_{d}$). By definition of $\Delta_{L}$, this has $q=x$ and $q=x_{d}$ as its particular solutions (see Example 5.1). As defined in the proof of Theorems 3.1, we have $d=\mu d_{c}\text{ and }G=\mu G_{c}$ for these virtual systems, resulting in the upper bounds of disturbances in the learning-based control formulation (5.18) and (5.22) given as
@@
Since $M$ and $u^{*}$ are constructed to render $\dot{q}=\zeta(q,\varpi,t)$ contracting as presented in Theorem 4.2, and we have $\|\Delta_{L}\|\leq\epsilon_{\ell 0}+\epsilon_{\ell 1}\|\xi_{1}-\xi_{0}\|$ due to the learning assumption (5.3), Theorem 5.2 implies the deterministic bound (6.9) and Theorem 5.3 implies the stochastic bounds (6.10) and (6.11). Note that using the differential feedback control of Theorem 4.6 as $u^{*}$ also gives the bounds (6.9)–(6.11) following the same argument Tsukamoto et al. [2021]. $\qed$
Due to the estimation and control duality in differential dynamics observed in Sec. 4, we have a similar result to Theorem 6.1 for the NCM-based state estimator.
@@
Let $\mathcal{M}$ define the NCM in Definition 6.2, and let $L(\hat{x},t)=M\bar{C}^{\top}R^{-1}$ for $M$, $\bar{C}$ and $R$ given in Theorem 4.5. Suppose that the systems (6.4) and (6.5) are estimated with an estimator gain $L_{L}$, which computes $L$ of the CV-STEM estimator $\dot{\hat{x}}=f(\hat{x},t)+L(\hat{x},t)(y-h(\hat{x},t))$ in Theorem 4.5 replacing $M$ by $\mathcal{M}$, i.e.,
@@
Define $\Delta_{L}$ of Theorems 5.2 and 5.3 as follows:
@@
where we use $q(0,t)=\xi_{0}(t)=x(t)$, $q(1,t)=\xi_{1}(t)=\hat{x}(t)$, and $\varpi=[x^{\top},\hat{x}^{\top}]^{\top}$ in the learning-based control formulation (5.1) and (5.2). If the NCM is learned to satisfy the learning error assumptions of Theorems 5.2 and 5.3 for $\Delta_{L}$ given by (6.15), then (5.18), (5.22), and (5.23) of Theorems 5.2 and 5.3 hold as follows:
@@
where $\mathtt{e}=\hat{x}-x$, $\bar{d}_{a}=\epsilon_{\ell 0}+\bar{d}_{e0}$, $\bar{d}_{b}=\bar{\rho}\bar{c}\bar{d}_{e1}$, $C_{E}=(\bar{g}_{e0}^{2}+\bar{\rho}^{2}\bar{c}^{2}\bar{g}_{e1}^{2}\overline{m}^{2})({2}{\alpha_{G}}^{-1}+1)+\epsilon_{\ell 0}^{2}\alpha_{d}^{-1}$, the disturbance upper bounds $\bar{d}_{e0}$ $\bar{d}_{e1}$, $\bar{g}_{e0}$, and $\bar{g}_{e1}$ are given in (6.4) and (6.5), and the other variables are defined in Theorems 4.55.2 and 5.3.
@@
Let g of the virtual systems given in Theorems 5.2 and 5.3 be $\textsl{g}=\zeta$, where $\zeta$ is given by (4.10) (i.e., $\zeta=(A-LC)(q-x)+f(x,t)$). By definition of $\Delta_{L}$, this has $q=\hat{x}$ and $q=x$ as its particular solutions (see Example 5.2). As defined in the proof Theorem  4.3, we have
@@
Since $W=M^{-1}$ is constructed to render $\dot{q}=\zeta(q,\varpi,t)$ contracting as presented in Theorem 4.5, and we have $\|\Delta_{L}\|\leq\epsilon_{\ell 0}+\epsilon_{\ell 1}\|\xi_{1}-\xi_{0}\|$ due to the learning assumption (5.3), Theorem 5.2 implies the bound (6.16), and Theorem 5.3 implies the bounds (6.17) and (6.18). $\qed$
@@
As discussed in Examples 5.1 and 5.2 and in Remark 5.1, we have $\epsilon_{\ell 1}=0$ for the learning error $\epsilon_{\ell 1}$ of (5.3) as long as $B$ and $h$ are bounded a compact set by the definition of $\Delta_{L}$ in (6.8) and (6.15). If $h$ and $u^{*}$ are Lipschitz with respect to $x$ in the compact set, we have non-zero $\epsilon_{\ell 1}$ with $\epsilon_{\ell 0}=0$ in (5.3) (see Theorem 6.3).
Using Theorems 6.1 and 6.2, we can relate the learning error of the matrix that defines the NCM, $\|\mathcal{M}-M\|$, to the error bounds (6.9)–(6.11) and (6.16)–(6.18), if $M$ is given by the CV-STEM of Theorems 4.2 and 4.5.
@@
Let $\mathcal{M}$ define the NCM in Definition 6.2, and let $\mathcal{S}_{s}\subseteq\mathbb{R}^{n}$, $\mathcal{S}_{u}\subseteq\mathbb{R}^{m}$, and $t\in\mathcal{S}_{t}\subseteq\mathbb{R}_{\geq 0}$ be some compact sets. Suppose that the systems (6.1)–(6.5) are controlled by (6.7) and estimated by (6.13), respectively, as in Theorems 6.1 and 6.2. Suppose also that $\exists\bar{b},\bar{c},\bar{\rho}\in\mathbb{R}_{\geq 0}$ s.t. $\|B(x,t)\|\leq\bar{b}$, $\|C(\varrho_{c},x,\hat{x},t)\|\leq\bar{c}$, and $\|R^{-1}\|\leq\bar{\rho},~\forall x,\hat{x}\in\mathcal{S}_{s}$ and $t\in\mathcal{S}_{t}$, for $B$ in (6.1), $C$ in Theorem 4.5, and $R$ in (6.7) and (6.13). If we have for all $x,x_{d},\hat{x}\in\mathcal{S}_{s}$, $u_{d}\in\mathcal{S}_{u}$, and $t\in\mathcal{S}_{t}$ that
@@
where $M(x,x_{d},u_{d},t)$ and $M(\hat{x},t)$ (or $M(x,t)$ and $M(\hat{x},t)$, see Theorem 3.2) are the CV-STEM solutions of Theorems 4.2 and 4.5 with $\exists\epsilon_{\ell}\in\mathbb{R}_{\geq 0}$ being the learning error, then the bounds (6.9)–(6.11) of Theorem 6.1 and (6.16)–(6.18) of Theorem 6.2 hold with $\epsilon_{\ell 0}=0$ and $\epsilon_{\ell 1}=\bar{\rho}\bar{b}^{2}\epsilon_{\ell}$ for control, and $\epsilon_{\ell 0}=0$ and $\epsilon_{\ell 1}=\bar{\rho}\bar{c}^{2}\epsilon_{\ell}$ for estimation, as long as $\epsilon_{\ell}$ is sufficiently small to satisfy the conditions (5.17) and (5.21) of Theorems 5.2 and 5.3 for deterministic and stochastic systems, respectively.
@@
The left-hand side of Fig. 4 shows the state estimation errors of the following Lorenz oscillator perturbed by process noise $d_{0}$ and measurement noise $d_{1}$ Tsukamoto and Chung [2021]:
@@
where $x=[x_{1},x_{2},x_{3}]^{\top}$, $\sigma=10$, $\beta=8/3$, and $\rho=28$. It can be seen that the steady-state estimation errors of the NCM and CV-STEM are below the optimal steady-state upper bound of Theorem 4.5 (dotted line in the left-hand side of Fig. 4) while the EKF has a larger error compared to the other two. In this example, training data is sampled along $100$ trajectories with uniformly distributed initial conditions ($10\leq x_{i}\leq 10,~i=1,2,3$) using Theorem 4.5, and then modeled by a DNN as in Definition 6.2. Additional implementation and network training details can be found in Tsukamoto and Chung [2021].
@@
Consider a robust feedback control problem of the planar spacecraft perturbed by deterministic disturbances, the unperturbed dynamics of which is given as follows Nakka et al. [2018]:
@@
where $x=[p_{x},p_{y},\phi,\dot{p}_{x},\dot{p}_{y},\dot{\phi}]^{T}$ with $p_{x}$, $p_{y}$, and $\phi$ being the horizontal coordinate, vertical coordinate, and yaw angle of the spacecraft, respectively, with the other variables defined as $m=\text{mass of spacecraft}$, $I=\text{yaw moment of inertia}$, $u=\text{thruster force vector}$, $\ell=\text{half-depth of spacecraft}$, and $b=\text{half-width of spacecraft}$ (see Fig. 8 of Nakka et al. [2018] for details).
As shown in the right-hand side of Fig. 4, the NCM keeps their trajectories within the bounded error tube of Theorem 4.2 (shaded region) around the target trajectory (dotted line) avoiding collision with the circular obstacles, even under the presence of disturbances, whilst requiring much smaller computational cost than the CV-STEM. Data sampling and the NCM training are performed as in Example 6.1 Tsukamoto and Chung [2021].
Figure 4: Lorenz oscillator state estimation error smoothed using a $15$-point moving average filter in Example 6.1 (left), and spacecraft motion $(p_{x},p_{y})$ on a planar field in Example 6.2 (right).
@@
We could also simultaneously synthesize a controller $u$ and the NCM, and Theorem 6.1 can provide formal robustness and stability guarantees even for such cases Sun et al. [2020], Tsukamoto et al. [2021]. In Sec. 7, we generalize the NCM to learn contraction theory-based robust control using a DNN only taking $x$, $t$, and a vector containing local environment information as its inputs, to avoid the online computation of $x_{d}$ for the sake of automatic guidance and control implementable in real-time.
6.2 NCMs as Lyapunov Functions
We could also utilize the NCM constructed by the CV-STEM contraction metric for designing a Control Lyapunov function (CLF) Sontag [1983, 1989], Khalil [2002], resulting in a stability result different from Theorems 6.1 and 6.3. To this end, let us recall that designing optimal $k$ of $u=k(x,x_{d},u_{d},t)$ reduces to designing optimal $K(x,x_{d},u_{d},t)$ of $u=u_{d}-K(x,x_{d},u_{d},t)(x-x_{d})$ due to Lemma 3.2 of Sec. 3. For such $u$, the virtual system of (6.1)/(6.2) and (6.3) without any perturbation, which has $q=x$ and $q=x_{d}$ as its particular solutions, is given as follows:

### v8

@@ 5.3. Motivation for Neural Contraction Metrics
The CV-STEM schemes in Theorems $LABEL:THM:Thm:CV-STEM$-estimation, and LABEL:THM:Thm:ccm_cvstem permit the construction of optimal contraction metrics via convex optimization for synthesizing optimal, provably stable, and robust feedback control and state estimation. It is also shown in Theorems LABEL:THM:Thm:contraction_learning and LABEL:THM:Thm:contraction_learning_sto that these contraction metrics are useful for obtaining stability guarantees for machine learning-based and data-driven control techniques while retaining the CV-STEM-type optimality due to the tracking error bounds (5.18) and (5.22), which are comparable to those in Theorems LABEL:THM:Thm:CVSTEMobjective and LABEL:THM:Thm:CV-STEMobjective-est. However, the CV-STEM requires that a nonlinear system of equations or an optimization problem be solved at each time instant, which is not suitable for systems with limited online computational capacity.
@@
The Neural Contraction Metric (NCM) and its extensions ncm, nscm, ancm, lagros have been developed to address such a computational issue by modeling the CV-STEM optimization scheme using a DNN neural1, neural2, neural3, making it implementable in real-time. We will see in the subsequent sections that Theorems LABEL:THM:Thm:contraction_learning and LABEL:THM:Thm:contraction_learning_sto provide one formal approach to prove robustness and stability properties of the NCM methodologies cdc_ncm.
6. Learning-based Robust Control and Estimation
@@
with $\sup_{x,t}\|d_{e0}\|=\bar{d}_{e0}\in\mathbb{R}_{\geq 0}$, $\sup_{x,t}\|d_{e1}\|=\bar{d}_{e1}\in\mathbb{R}_{\geq 0}$, $\sup_{x,t}\|G_{e0}\|_{F}=\bar{g}_{e0}\in\mathbb{R}_{\geq 0}$, and $\sup_{x,t}\|G_{e1}\|_{F}=\bar{g}_{e1}\in\mathbb{R}_{\geq 0}$. The following definitions of a DNN neural1, neural2, neural3 and NCM ncm, nscm, cdc_ncm will be utilized extensively hereinafter.
@@
Neural NetworksDef:NN_definition A neural network is a nonlinear mathematical model to approximately represent training samples $\{(x_{i},y_{i})\}_{i=1}^{N}$ of $y=\phi(x)$ by optimally tuning its hyperparameters $W_{\ell}$, and is given as follows:
@@
Neural Contraction MetricsDef:NCM A Neural Contraction Metric (NCM) is a DNN model of a contraction metric given in Theorem LABEL:THM:Thm:contraction. Its training data could be sampled by solving the CV-STEM presented in Theorem LABEL:THM:Thm:CV-STEM for control (or Theorem LABEL:THM:Thm:ccm_cvstem for differential feedback control) and Theorem LABEL:THM:Thm:CV-STEM-estimation for estimation. The NCM framework is summarized in Fig. 3.
@@
Although a stochastic version of the NCM is called a Neural Stochastic Contraction Metric (NSCM) in nscm, we also denote it as an NCM in this paper for simplicity of presentation.
@@
6.1. Stability of NCM-based Control and Estimation
@@
Learning-based Control with Neural Contraction MetricsThm:NCMstability1 Let $\mathcal{M}$ define the NCM in Definition LABEL:DEF:Def:NCM, and let $u^{*}=u_{d}-R^{-1}B^{\top}M(x-x_{d})$ for $M$ and $R$ given in Theorem LABEL:THM:Thm:CV-STEM. Suppose that the systems (6.1) and (6.2) are controlled by $u_{L}$, which computes the CV-STEM controller $u^{*}$ of Theorem LABEL:THM:Thm:CV-STEM replacing $M$ by $\mathcal{M}$, i.e.,
@@
for $\mathtt{e}=x-x_{d}$. Note that we could use the differential feedback control $u=u_{d}+\int^{1}_{0}kd\mu$ of Theorem LABEL:THM:Thm:ccm_cvstem for $u^{*}$ and $u_{L}$. Define $\Delta_{L}$ of Theorems LABEL:THM:Thm:contraction_learning and LABEL:THM:Thm:contraction_learning_sto as follows:
@@
where we use $q(0,t)=\xi_{0}(t)=x_{d}(t)$, $q(1,t)=\xi_{1}(t)=x(t)$, and $\varpi=[x^{\top},x_{d}^{\top},u_{d}^{\top}]^{\top}$ in the learning-based control formulation (5.1) and (5.2). If the NCM is learned to satisfy the learning error assumptions of Theorems LABEL:THM:Thm:contraction_learning and LABEL:THM:Thm:contraction_learning_sto for $\Delta_{L}$ given by (6.8), then (5.18), (5.22), and (5.23) of Theorems LABEL:THM:Thm:contraction_learning and LABEL:THM:Thm:contraction_learning_sto hold as follows:
@@
where $\mathtt{e}=x-x_{d}$, $C_{C}=\bar{g}_{c}^{2}({2}{\alpha_{G}}^{-1}+1)+\epsilon_{\ell 0}^{2}\alpha_{d}^{-1}$, the disturbance upper bounds $\bar{d}_{c}$ and $\bar{g}_{c}$ are given in (6.1) and (6.2), respectively, and the other variables are defined in Theorems LABEL:THM:Thm:CV-STEMLABEL:THM:Thm:contraction_learning and LABEL:THM:Thm:contraction_learning_sto.
@@
Let g of the virtual systems given in Theorems LABEL:THM:Thm:contraction_learning and LABEL:THM:Thm:contraction_learning_sto be $\textsl{g}=\zeta$, where $\zeta$ is given by (3.19) (i.e., $\zeta=(A-BR^{-1}B^{\top}M)(q-x_{d})+\dot{x}_{d}$). By definition of $\Delta_{L}$, this has $q=x$ and $q=x_{d}$ as its particular solutions (see Example LABEL:EX:ex:learning_prob1). As defined in the proof of Theorems LABEL:THM:Thm:CVSTEM:LMI, we have $d=\mu d_{c}\text{ and }G=\mu G_{c}$ for these virtual systems, resulting in the upper bounds of disturbances in the learning-based control formulation (5.18) and (5.22) given as
@@
Since $M$ and $u^{*}$ are constructed to render $\dot{q}=\zeta(q,\varpi,t)$ contracting as presented in Theorem LABEL:THM:Thm:CV-STEM, and we have $\|\Delta_{L}\|\leq\epsilon_{\ell 0}+\epsilon_{\ell 1}\|\xi_{1}-\xi_{0}\|$ due to the learning assumption (5.3), Theorem LABEL:THM:Thm:contraction_learning implies the deterministic bound (6.9) and Theorem LABEL:THM:Thm:contraction_learning_sto implies the stochastic bounds (6.10) and (6.11). Note that using the differential feedback control of Theorem LABEL:THM:Thm:ccm_cvstem as $u^{*}$ also gives the bounds (6.9)–(6.11) following the same argument cdc_ncm. $\qed$
Due to the estimation and control duality in differential dynamics observed in Sec. 4, we have a similar result to Theorem LABEL:THM:Thm:NCMstability1 for the NCM-based state estimator.
@@
Learning-based Estimation with Neural Contraction MetricsThm:NCMstability1_est Let $\mathcal{M}$ define the NCM in Definition LABEL:DEF:Def:NCM, and let $L(\hat{x},t)=M\bar{C}^{\top}R^{-1}$ for $M$, $\bar{C}$ and $R$ given in Theorem LABEL:THM:Thm:CV-STEM-estimation. Suppose that the systems (6.4) and (6.5) are estimated with an estimator gain $L_{L}$, which computes $L$ of the CV-STEM estimator $\dot{\hat{x}}=f(\hat{x},t)+L(\hat{x},t)(y-h(\hat{x},t))$ in Theorem LABEL:THM:Thm:CV-STEM-estimation replacing $M$ by $\mathcal{M}$, i.e.,
@@
Define $\Delta_{L}$ of Theorems LABEL:THM:Thm:contraction_learning and LABEL:THM:Thm:contraction_learning_sto as follows:
@@
where we use $q(0,t)=\xi_{0}(t)=x(t)$, $q(1,t)=\xi_{1}(t)=\hat{x}(t)$, and $\varpi=[x^{\top},\hat{x}^{\top}]^{\top}$ in the learning-based control formulation (5.1) and (5.2). If the NCM is learned to satisfy the learning error assumptions of Theorems LABEL:THM:Thm:contraction_learning and LABEL:THM:Thm:contraction_learning_sto for $\Delta_{L}$ given by (6.15), then (5.18), (5.22), and (5.23) of Theorems LABEL:THM:Thm:contraction_learning and LABEL:THM:Thm:contraction_learning_sto hold as follows:
@@
where $\mathtt{e}=\hat{x}-x$, $\bar{d}_{a}=\epsilon_{\ell 0}+\bar{d}_{e0}$, $\bar{d}_{b}=\bar{\rho}\bar{c}\bar{d}_{e1}$, $C_{E}=(\bar{g}_{e0}^{2}+\bar{\rho}^{2}\bar{c}^{2}\bar{g}_{e1}^{2}\overline{m}^{2})({2}{\alpha_{G}}^{-1}+1)+\epsilon_{\ell 0}^{2}\alpha_{d}^{-1}$, the disturbance upper bounds $\bar{d}_{e0}$ $\bar{d}_{e1}$, $\bar{g}_{e0}$, and $\bar{g}_{e1}$ are given in (6.4) and (6.5), and the other variables are defined in Theorems LABEL:THM:Thm:CV-STEM-estimationLABEL:THM:Thm:contraction_learning and LABEL:THM:Thm:contraction_learning_sto.
@@
Let g of the virtual systems given in Theorems LABEL:THM:Thm:contraction_learning and LABEL:THM:Thm:contraction_learning_sto be $\textsl{g}=\zeta$, where $\zeta$ is given by (4.10) (i.e., $\zeta=(A-LC)(q-x)+f(x,t)$). By definition of $\Delta_{L}$, this has $q=\hat{x}$ and $q=x$ as its particular solutions (see Example LABEL:EX:ex:learning_prob2). As defined in the proof Theorem  LABEL:THM:Thm:CVSTEM-LMI-est, we have
@@
Since $W=M^{-1}$ is constructed to render $\dot{q}=\zeta(q,\varpi,t)$ contracting as presented in Theorem LABEL:THM:Thm:CV-STEM-estimation, and we have $\|\Delta_{L}\|\leq\epsilon_{\ell 0}+\epsilon_{\ell 1}\|\xi_{1}-\xi_{0}\|$ due to the learning assumption (5.3), Theorem LABEL:THM:Thm:contraction_learning implies the bound (6.16), and Theorem LABEL:THM:Thm:contraction_learning_sto implies the bounds (6.17) and (6.18). $\qed$
@@
As discussed in Examples LABEL:EX:ex:learning_prob1 and LABEL:EX:ex:learning_prob2 and in Remark 5.1, we have $\epsilon_{\ell 1}=0$ for the learning error $\epsilon_{\ell 1}$ of (5.3) as long as $B$ and $h$ are bounded a compact set by the definition of $\Delta_{L}$ in (6.8) and (6.15). If $h$ and $u^{*}$ are Lipschitz with respect to $x$ in the compact set, we have non-zero $\epsilon_{\ell 1}$ with $\epsilon_{\ell 0}=0$ in (5.3) (see Theorem LABEL:THM:Thm:NCMstability2).
Using Theorems LABEL:THM:Thm:NCMstability1 and LABEL:THM:Thm:NCMstability1_est, we can relate the learning error of the matrix that defines the NCM, $\|\mathcal{M}-M\|$, to the error bounds (6.9)–(6.11) and (6.16)–(6.18), if $M$ is given by the CV-STEM of Theorems LABEL:THM:Thm:CV-STEM and LABEL:THM:Thm:CV-STEM-estimation.
@@
Robustness against Metric Construction ErrorsThm:NCMstability2 Let $\mathcal{M}$ define the NCM in Definition LABEL:DEF:Def:NCM, and let $\mathcal{S}_{s}\subseteq\mathbb{R}^{n}$, $\mathcal{S}_{u}\subseteq\mathbb{R}^{m}$, and $t\in\mathcal{S}_{t}\subseteq\mathbb{R}_{\geq 0}$ be some compact sets. Suppose that the systems (6.1)–(6.5) are controlled by (6.7) and estimated by (6.13), respectively, as in Theorems LABEL:THM:Thm:NCMstability1 and LABEL:THM:Thm:NCMstability1_est. Suppose also that $\exists\bar{b},\bar{c},\bar{\rho}\in\mathbb{R}_{\geq 0}$ s.t. $\|B(x,t)\|\leq\bar{b}$, $\|C(\varrho_{c},x,\hat{x},t)\|\leq\bar{c}$, and $\|R^{-1}\|\leq\bar{\rho},~\forall x,\hat{x}\in\mathcal{S}_{s}$ and $t\in\mathcal{S}_{t}$, for $B$ in (6.1), $C$ in Theorem LABEL:THM:Thm:CV-STEM-estimation, and $R$ in (6.7) and (6.13). If we have for all $x,x_{d},\hat{x}\in\mathcal{S}_{s}$, $u_{d}\in\mathcal{S}_{u}$, and $t\in\mathcal{S}_{t}$ that
@@
where $M(x,x_{d},u_{d},t)$ and $M(\hat{x},t)$ (or $M(x,t)$ and $M(\hat{x},t)$, see Theorem LABEL:THM:Thm:fixed_sdc) are the CV-STEM solutions of Theorems LABEL:THM:Thm:CV-STEM and LABEL:THM:Thm:CV-STEM-estimation with $\exists\epsilon_{\ell}\in\mathbb{R}_{\geq 0}$ being the learning error, then the bounds (6.9)–(6.11) of Theorem LABEL:THM:Thm:NCMstability1 and (6.16)–(6.18) of Theorem LABEL:THM:Thm:NCMstability1_est hold with $\epsilon_{\ell 0}=0$ and $\epsilon_{\ell 1}=\bar{\rho}\bar{b}^{2}\epsilon_{\ell}$ for control, and $\epsilon_{\ell 0}=0$ and $\epsilon_{\ell 1}=\bar{\rho}\bar{c}^{2}\epsilon_{\ell}$ for estimation, as long as $\epsilon_{\ell}$ is sufficiently small to satisfy the conditions (5.17) and (5.21) of Theorems LABEL:THM:Thm:contraction_learning and LABEL:THM:Thm:contraction_learning_sto for deterministic and stochastic systems, respectively.
@@
Lorenz Oscillatorex:lorenz The left-hand side of Fig. 4 shows the state estimation errors of the following Lorenz oscillator perturbed by process noise $d_{0}$ and measurement noise $d_{1}$ ncm:
@@
where $x=[x_{1},x_{2},x_{3}]^{\top}$, $\sigma=10$, $\beta=8/3$, and $\rho=28$. It can be seen that the steady-state estimation errors of the NCM and CV-STEM are below the optimal steady-state upper bound of Theorem LABEL:THM:Thm:CV-STEM-estimation (dotted line in the left-hand side of Fig. 4) while the EKF has a larger error compared to the other two. In this example, training data is sampled along $100$ trajectories with uniformly distributed initial conditions ($10\leq x_{i}\leq 10,~i=1,2,3$) using Theorem LABEL:THM:Thm:CV-STEM-estimation, and then modeled by a DNN as in Definition LABEL:DEF:Def:NCM. Additional implementation and network training details can be found in ncm.
@@
Planar Spacecraft Simulatorsex:sc_robust_control Consider a robust feedback control problem of the planar spacecraft perturbed by deterministic disturbances, the unperturbed dynamics of which is given as follows SCsimulator:
@@
where $x=[p_{x},p_{y},\phi,\dot{p}_{x},\dot{p}_{y},\dot{\phi}]^{T}$ with $p_{x}$, $p_{y}$, and $\phi$ being the horizontal coordinate, vertical coordinate, and yaw angle of the spacecraft, respectively, with the other variables defined as $m=\text{mass of spacecraft}$, $I=\text{yaw moment of inertia}$, $u=\text{thruster force vector}$, $\ell=\text{half-depth of spacecraft}$, and $b=\text{half-width of spacecraft}$ (see Fig. 8 of SCsimulator for details).
As shown in the right-hand side of Fig. 4, the NCM keeps their trajectories within the bounded error tube of Theorem LABEL:THM:Thm:CV-STEM (shaded region) around the target trajectory (dotted line) avoiding collision with the circular obstacles, even under the presence of disturbances, whilst requiring much smaller computational cost than the CV-STEM. Data sampling and the NCM training are performed as in Example LABEL:EX:ex:lorenz ncm.
Figure 4: Lorenz oscillator state estimation error smoothed using a $15$-point moving average filter in Example LABEL:EX:ex:lorenz (left), and spacecraft motion $(p_{x},p_{y})$ on a planar field in Example LABEL:EX:ex:sc_robust_control (right).
@@
We could also simultaneously synthesize a controller $u$ and the NCM, and Theorem LABEL:THM:Thm:NCMstability1 can provide formal robustness and stability guarantees even for such cases chuchu, cdc_ncm. In Sec. 7, we generalize the NCM to learn contraction theory-based robust control using a DNN only taking $x$, $t$, and a vector containing local environment information as its inputs, to avoid the online computation of $x_{d}$ for the sake of automatic guidance and control implementable in real-time.
6.2. NCMs as Lyapunov Functions
We could also utilize the NCM constructed by the CV-STEM contraction metric for designing a Control Lyapunov function (CLF) doi:10.1137/0321028, SONTAG1989117, Khalil:1173048, resulting in a stability result different from Theorems LABEL:THM:Thm:NCMstability1 and LABEL:THM:Thm:NCMstability2. To this end, let us recall that designing optimal $k$ of $u=k(x,x_{d},u_{d},t)$ reduces to designing optimal $K(x,x_{d},u_{d},t)$ of $u=u_{d}-K(x,x_{d},u_{d},t)(x-x_{d})$ due to Lemma 3.2 of Sec. 3. For such $u$, the virtual system of (6.1)/(6.2) and (6.3) without any perturbation, which has $q=x$ and $q=x_{d}$ as its particular solutions, is given as follows: